Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 12,000 $ 16,200 $ 28,200
Estimated variable manufacturing overhead per machine-hour $ 2.20 $ 3.00
Job P Job Q
Direct materials $ 21,000 $ 12,000
Direct labor cost $ 27,400 $ 10,700
Actual machine-hours used:
Molding 2,500 1,600
Fabrication 1,400 1,700
Total 3,900 3,300

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

12.If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations.)

- If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

-Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

-What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)
Answer #1
Molding Fabrication
Estimated total fixed manufacturing overhead 12000 16200
Estimated variable manufacturing overhead 5500 4500
Total manufacturing overhead 17500 20700
Divide by Total machine-hours used 2500 1500
Departmental predetermined overhead rates 7.00 13.80
12
Job P
Direct materials 21000
Direct labor cost 27400
Overhead applied:
Molding 17500 =2500*7
Fabrication 19320 =1400*13.8
Total cost 85220
Divide by units 20
Unit product cost Job P 4261
13
Job Q
Direct materials 12000
Direct labor cost 10700
Overhead applied:
Molding 11200 =1600*7
Fabrication 23460 =1700*13.8
Total cost 57360
Divide by units 30
Unit product cost Job Q 1912
14
Job P Job Q
Total cost 85220 57360
Add: Markup @ 80% 68176 45888
Total selling price for the job 153396 103248
Divide by units 20 30
Selling price per unit 7670 3442
15
Total cost Job P 85220
Total cost Job Q 57360
Cost of goods sold for March 142580